# Using Reinforcement Learning to Train Large Language Models to Explain Human Decisions

### v1

@@ E.2 Cognitive mechanisms identified by Gemma-2-2B-Instruct
Appendix F Implementation details
7B models. RL training was conducted over 3 epochs using the training set, distributed across 4$\times$×H100 GPUs for a total runtime of 80 hours. For SFT and Centaur-style SFT trainings, each model was trained for 6 epochs on the training set, using a single A100 GPU with an approximate runtime of 5 hours per training session.
Inference across all checkpoints on the complete choices13k dataset took approximately 5 hours on a single A100 GPU for the SFT and Centaur-style SFT models using vLLM [19], whereas the RL models required roughly 30 hours under the same conditions.
2B models. RL training was conducted over 3 epochs using the training set, distributed across 4$\times$×H100 GPUs for a total runtime of 40 hours. For SFT and Centaur-style SFT trainings, each model was trained for 6 epochs on the training set, using a single A100 GPU with an approximate runtime of 3 hours per training session.
Inference across all checkpoints on the complete choices13k dataset took approximately 3 hours on a single A100 GPU for the SFT and Centaur-style SFT models using vLLM [19], whereas the RL models required roughly 20 hours under the same conditions.
Preliminary or failed experiments. We also conducted several pilot studies and unsuccessful RL runs, which together accounted for approximately 2,000 H100 GPU hours.

### v2

@@ E.2 Cognitive mechanisms identified by Gemma-2-2B-Instruct
Appendix F Additional discussion
Failures of RL. During our experiments, we also observed several failures associated with GRPO training. Most notably, applying the original GRPO algorithm (Shao et al., 2024) to the choices13k dataset led to mode collapse. By mode collapse, we refer to the phenomenon in which the fine-tuned LLMs converge to generating a single, repetitive reasoning chain that simply computes the expected values of both options and consistently outputs a final prediction of 50% for each option across all choice problems. Note that the choices13k dataset exhibits a modal human choice proportion of 50%, which may have contributed to this degeneracy.
We experimented with several strategies to prevent the model from collapsing to this default prediction. Specifically, we tested alternative reward functions, including one minus the mean squared error, negative cross-entropy loss, and the addition of diversity bonuses incentivizing diverse predictions. However, none of these modifications were effective in mitigating mode collapse in our setting. The issue was eventually addressed by removing advantage normalization in GRPO, as recommended by Liu et al. (2025). That is, dividing the centered reward by the standard deviation of group rewards introduces a bias toward question-level difficulty: risky-choice problems with lower standard deviations (i.e., those that are either too easy or too difficult) disproportionately influence policy updates by receiving higher weights. By omitting this normalization step, we were able to recover the main RL results reported above.
After removing normalization, the stability of RL training improved substantially. We then evaluated three reward functions under GRPO without normalization: (i) $1-\mbox{absolute prediction error}$, (ii) $1-\mbox{squared error}$, and (iii) negative cross-entropy loss. Among these, reward function (i) yielded the most stable RL training, characterized by smoother KL trajectories and smaller reward standard deviations compared to (ii) and (iii). We suspect this advantage arises because (i) aligns more closely with GRPO’s centered reward design.
The elicitation hypothesis. A prominent hypothesis about RL post-training is that it primarily increases the probability of generating correct outputs by eliciting behaviors already latent in the pretrained LLM (Lambert, 2025; Yue et al., 2025). In this view, RL post-training does not teach new capabilities but instead amplifies and selects from pre-existing knowledge. Our model control results support this hypothesis: weaker LLMs — particularly those unable to elicit the expected value mechanism for risky choice — struggle to explain human behavior effectively when trained with RL. These results suggest that the success of RL highly depends on the capacity of the backbone LLM.
Appendix G Implementation details
[…]
Inference across all checkpoints on the complete choices13k dataset took approximately 5 hours on a single A100 GPU for the SFT and Centaur-style SFT models using vLLM (Kwon et al., 2023), whereas the RL models required roughly 30 hours under the same conditions.
[…]
Inference across all checkpoints on the complete choices13k dataset took approximately 3 hours on a single A100 GPU for the SFT and Centaur-style SFT models using vLLM (Kwon et al., 2023), whereas the RL models required roughly 20 hours under the same conditions.
Pilot and control experiments. We also conducted several RL runs for pilot and control experiments, which together accounted for approximately 2,500 H100 GPU hours.